The following information applies to the questions displayed below.]

Clarke Corporation manufactures three products (X-1, X-2, and X-3) utilizing, in one of the processes, a single machine for all products. Data on the individual products follow.

X-1 X-2 X-3
Price per unit $ 160 $ 260 $ 520
Variable cost per unit $ 80 $ 140 $ 300
Machine hours per unit 1.0 2.5 4.0
Maximum units demand per period 375,000 180,000 65,000

The single machine used for all three products has a maximum capacity of 798,000 hours per period.

Required:

a. How many units of each product should be produced each period?

b. A local engineering firm has suggested to Clarke that it might be able to increase the capacity on the machine. What is the maximum Clarke would be willing to pay for an increase of twenty (20) machine hours?

c. Suppose the maximum capacity on the machine was 435,000 hours (instead of 798,000 hours). What is the maximum Clarke would be willing to pay for an increase of twenty (20) machine hours?

d. At what capacity (in machine hours) would the machine no longer be a bottleneck?

solve all 4 parts asap thanks
Answer #1
Calculation of contribution margin per machine hour
X-1 X-2 X-3
Price per unit 160 260 520
Variable cost per unit 80 140 300
Contribution Margin per Unit 80 120 220
Machine hours per unit 1 2.5 4
Contribution Margin per Machine hour 80 48 55
Production Schedule
Product Number of Units Machine Hours per Unit Total Machine Hours
X-1 375,000 1 375,000
X-2 65,200 2.5 163,000
X-3 65,000 4 260,000
Total 798,000
Maximum for increase = contribution margin earned = 20*48 = $960
c.Increased machine hours will be used to produce X-3, hence amount paid = 20*55 = $1100
d.Capacity in machine hours = 375,000+260,000+180,000*2.5 = 1,085,000 hours
i.e. when demand for all products will be met